### v1

# Multi-color characterization of optically invisible FU Orionis-type outbursts: Demonstration and prospects for the WINTER survey

Abstract
Episodic mass accretion is the dominant mechanism for mass assembly in the proto-stellar phase. Although prior optical time-domain searches have allowed detailed studies of individual outbursts, these searches remain insensitive to the earliest stages of star formation. In this paper, we present the characterization of two FU Orionis (FUor) outbursts identified using the combination of the ground-based, near-infrared Wide-field Infrared Transient Explorer (WINTER) and the space-based, mid-infrared NEOWISE survey. Supplemented with near-infrared spectroscopic follow-up, we show that both objects are bona fide FUor type outbursts based on i) their proximity to star-forming regions, ii) large amplitude (2-4 magnitudes) infrared brightening over the last decade, iii) progenitor colors consistent with embedded (Class I) protostars, and iv) “mixed-temperature” infrared spectra exhibiting characteristic signatures of cool outer envelopes and a hot inner disk with a wind. While one source, WNTR24-cua, is a known FUor which we independently recover; the second source, WNTR24-egv, is a newly confirmed object. Neither source is detected in contemporaneous ground-based optical imaging, despite flux limits fainter than their infrared brightness, demonstrating the capabilities of WINTER to identify heavily obscured young stellar object (YSO) outbursts. We highlight the capabilities of the Galactic Plane survey of the recently commissioned WINTER observatory in addressing the poorly understood FUor population with its unique combination of real-time detection capabilities, multi-color sensitivity, weekly cadence, and wide area coverage.
1 Introduction
Young stellar objects (YSOs) grow through episodic accretion, where material from a surrounding disk falls onto the star in bursts rather than at a steady rate (Fischer et al., 2023, and references therein). These accretion outbursts significantly increase the system’s luminosity and play a major role in stellar mass assembly (Kenyon et al., 1990; Hartmann & Kenyon, 1996; Vorobyov & Basu, 2005). Episodic accretion has been proposed as a solution to the “luminosity problem,” in which protostars appear dimmer than expected if they were gaining mass at a constant rate throughout their evolution (Kenyon et al., 1990; Dunham & Vorobyov, 2012). In addition to influencing how stars gain mass, these bursts can alter the thermal and chemical structure of the disk, affecting processes such as snowline migration and early planet formation (Nayakshin et al., 2023; Lee et al., 2019; Jørgensen et al., 2013; Kóspál et al., 2023). Discovering and characterizing these outbursts remains essential for understanding their physical origins and role in star formation.
The most extreme form of episodic accretion in YSOs is FU Orionis-type (FUor) outbursts, characterized by large luminosity increases (mag 4–6) and lasting decades or more (Hartmann & Kenyon, 1996; Audard et al., 2014; Fischer et al., 2023). FUor outbursts sustain accretion rates of up to , contributing significantly to stellar mass growth in the earliest stages of evolution (Hartmann & Kenyon, 1996; Vorobyov & Basu, 2015). In contrast, another common category of bursts, EX Lupi-type outbursts (EXors), are less luminous (mag 2–3), shorter-lived (weeks to a few years), and contribute less significantly to the stellar mass budget (Ashraf et al., 2024; Cruz-Sáenz de Miera et al., 2023). Despite the proposed significance of FUors in mass accumulation, they are observationally rare. The classic compilation of Connelley & Reipurth (2018) reported fewer than 35 FUor-type events had been confirmed using spectroscopic techniques, albeit with most objects having sparse or no photometric coverage of the outburst.
YSO variability, including FUor outbursts, is commonly studied through time-domain surveys across optical and infrared (IR) wavelengths, often with follow-up spectroscopy for classification. Early discoveries and monitoring of eruptive YSOs relied heavily on optical time-domain surveys, with discoveries in surveys such as ASAS-SN (Herczeg et al., 2016), PTF (Miller et al., 2011), and, more recently, Gaia (Hillenbrand et al., 2018; Szegedi-Elek et al., 2020). However, optical time-domain surveys have limited sensitivity to embedded sources as well as regions of heavy foreground extinction in the Galactic plane and star-forming regions (Enoch et al., 2009; Dunham et al., 2015).
Recent wide-field IR surveys have proven essential for uncovering outbursts in heavily embedded YSOs, including classes of FUor outbursts previously missed in optical monitoring. The VISTA Variables in the Via Lactea (VVV; Minniti et al. 2010) and its extension VVVX enabled the discovery of dozens of high-amplitude eruptive YSOs, including several FUor candidates in the Galactic plane (Contreras Peña et al., 2017a, b, 2024; Guo et al., 2024a). Similarly, mid-infrared (MIR) monitoring with WISE (Wright et al., 2010) and NEOWISE (Mainzer et al., 2014) uncovered a growing population of deeply embedded eruptive variables (Hillenbrand et al., 2018; Ashraf et al., 2024; Tran et al., 2024; Contreras Peña et al., 2025). While the VVV and NEOWISE searches were largely implemented via archival photometric searches together with spectroscopic follow-up many years after the outburst onset, the advent of real-time transient discovery techniques is enabling prompt identification and early spectroscopic follow-up of these outbursts. The ongoing Palomar Gattini-IR (PGIR) survey has uncovered several new NIR-bright FUor events in the northern hemisphere (Hillenbrand et al., 2021, 2025) and continues to provide a key source of detections, especially toward the Galactic plane (Earley et al., in prep.).
In addition to discovery, multi-color infrared photometry and broad-band spectroscopy play a key role in confirming YSO outbursts and enabling detailed studies of accretion disk evolution. These observations also identify and exclude interlopers, including Mira variables, cataclysmic variables, and classical novae. Multi-wavelength monitoring probes accretion disk formation via changes in NIR and MIR colors—where delays between shorter and longer wavelength bands are critical to understanding the outburst triggering mechanism and instability propagation (outside-in or inside-out; Bell et al. 1995; Cleaver et al. 2023; Nayakshin & Elbakyan 2024). Simultaneously, such coverage traces the development of hot inner disk regions, the clearing or rebuilding of circumstellar material and the presence of infalling envelopes or outflows (Hillenbrand & Rodriguez, 2022; Caratti o Garatti et al., 2017). Connelley & Reipurth (2018) demonstrate that NIR spectroscopy provides a robust basis for classifying FUor-type objects, which exhibit a distinctive combination of strong molecular absorption, weak or absent emission lines, low-gravity spectral features, and a mixed-temperature spectrum that becomes progressively cooler at longer wavelengths.
The Wide-Field Infrared Transient Explorer (WINTER) is a new NIR time-domain survey on a dedicated 1-meter robotic telescope at Palomar Observatory (Lourie et al., 2020; Frostig et al., 2024). Each pointing covers a 1.2 field of view with a 1.1” pixel scale and the option of observing in Y, J, or a shortened H (Hs) filters, centered on 1.02, 1.25, and 1.60 . WINTER covers the Galactic plane during the summer months, running multiple surveys at an approximately two-week cadence. Unlike VVV and PGIR which almost entirely operated their time-domain surveys in a single filter (in Ks and J band, respectively), WINTER carries out multi-color cadenced surveys of the Galactic plane. Like PGIR, WINTER also includes the infrastructure for rapid discovery of transients with a real-time image subtraction pipeline, operating with smaller field of view (1.2 vs. 25 ) but substantially higher depth (median vs. AB mag) and spatial resolution (1.1″vs 8.7″) than PGIR (De et al., 2020).
In this study, we demonstrate WINTER’s potential for discovery and characterization of YSO outbursts via the detection and characterization of heavily obscured FUor outbursts identified independently via searches in NEOWISE. We present two FUors studied in NIR with WINTER, in MIR with WISE and NEOWISE, and with NIR spectroscopic follow up. One source, WNTR24-cua is a known YSO outburst, also known as 2017fzn and as L222_192 in Guo et al. (2024a), which also classified the object as an FUor. The second source, WNTR24-egv, is a newly characterized FUor, adding the expanding sample of optically obscured but IR-bright outbursts. Of note, WNTR24-egv was recently proposed as a candidate FUor by Contreras Peña et al. (2025), based on a long-term rise in NEOWISE photometry, although it has not been confirmed spectroscopically before this study.
We present the observations of WNTR24-cua and WNTR24-egv in WINTER, WISE, and other surveys, the source environment, and spectroscopic follow-up in Section 2. In Section 3, we analyze the observations, presenting a study of the host environments, the progenitors of the outbursts, and spectroscopically classify the FUors. We discuss the implications for future observations of YSO outbursts with WINTER and conclude in Section 4.
2 Observations
The two FUor events presented here were first identified as mid-infrared transients in systematic searches of NEOWISE data. As part of WINTER’s broader Galactic plane survey, we also targeted regions with NEOWISE outbursts to verify near-infrared brightness and obtain multi-color light curves. In this paper, we present detailed studies of two objects as a pilot demonstration of WINTER’s capabilities for characterizing obscured transients in the Galactic plane.
2.1 Discovery, NEOWISE, and WINTER initial observations
Systematic processing of NEOWISE data with the unWISE ZOGY pipeline (Lang, 2014; Meisner et al., 2018; De et al., 2023) flagged both sources as mid-infrared transients. The pipeline performs image subtraction between the original WISE coadds from 2010 and time-resolved NEOWISE coadds obtained from 2014–2024, producing a catalog of statistically significant transient candidates. The search identified both events as transients and assigned database names as part of the WISE Transient Pipeline (WTP). WNTR24-cua and WNTR24-egv are located at J2000 coordinates , and , and assigned WTP designations WTP20aabowt and WTP14acbxtp, respectively.
WINTER first observed both FUors in the band on UT 2024-07-10 (WNTR24-cua) and 2024-07-03 (WNTR24-egv) as part of the Galactic plane program. The data were processed with the WINTER pipeline implemented with mirar (Karambelkar and Stein et al., in prep), an open-source, modular python-based system. The pipeline runs all reduction steps including calibration, astrometry, stacking, difference imaging, and transient identification. Difference images were generated using reference images from UKIRT for and , and from Pan-STARRS , where available. Figure 1 shows example image cutouts produced by the pipeline from the original WINTER -band images, the UKIRT reference images, and the resulting difference images for both events. WNTR24-cua and WNTR24-egv were flagged as candidate transients and assigned WINTER database names, WNTR24gxcua and WNTR24fxegv. For readability throughout this work, we refer to them using shortened forms of these names.
The sources were selected for WINTER follow-up based on their high near-infrared brightness and coincidence with previously identified WISE transient candidates. Monitoring has continued through the 2025 Galactic plane season in the -, -, and -bands as part of WINTER’s ongoing infrared Galactic plane program.
2.2 Optical Observations
There are no optical detections of these IR-bright outbursts, with the exception of faint Gaia (Gaia Collaboration et al., 2016) -band coverage for WNTR24-cua. The Gaia -band has broad coverage from 111https://www.cosmos.esa.int/web/gaia/iow_20180316 , including some NIR flux around 1 . The Gaia data were downloaded from the Gaia Photometric Science Alerts webpage222https://gsaweb.ast.cam.ac.uk/alerts/, where the event is designated as Gaia 17bzk. These data include -band flux but no error bars. The Gaia alert data includes non-detections before the event and the decline of the outburst between UT 2017-08-07 and UT 2023-09-16 (as seen in Figure 2).
Archival data from the Zwicky Transient Facility (Bellm et al., 2019, ZTF), as accessed through the ZTF forced photometry server, revealed no detection of the outburst in available data between UT 2018-04-04 and UT 2024-12-20 in the more narrow , , or filters to an average 5 limiting depth of 20 mag AB in each band. This implies these sources are quite obscured at optical wavelengths, though bright at NIR and MIR wavelengths.
2.3 Light curve evolution
Both events have long-term infrared monitoring from the WISE and NEOWISE missions (Figure 2). They were first observed in two epochs during the original WISE mission in 2010, with continued coverage in the and bands by NEOWISE at a six-month cadence from 2014 through the final visit in 2024.
WNTR24-cua was not detected in WISE until early 2016, when it began a sharp rise in the band, brightening by 1.7 magnitudes over 364 days (Figure 2). This was followed by a decline of 0.5 magnitudes in 2019, after which the source exhibited a more gradual increase of 0.62 magnitudes between 2020 and 2024. The initial rise was not seen in , likely due to the band’s 1 magnitude shallower depth; however, near peak brightness the source was detected at mag, with a color of mag. By 2024, the emission became bluer with mag. The brightening is also evident in VVV , which shows the onset of the outburst in 2016, with the progenitor detected prior to the rise in , , and . In contrast, Gaia -band observations do not detect the source up through May 2017, but then observe the outburst three months later, implying a rapid optical rise after the start of the NIR and MIR outburst.
WNTR24-egv, in contrast, has been consistently detected in WISE since 2010. The earliest data show no significant variability, but a brightening of 0.9 (0.6) magnitudes in () began in 2014, following a brief dip, and continued steadily to an overall change of 3.7 (2.7) magnitudes on in () by 2024. The color has decreased throughout this evolution, beginning at 1.0 magnitudes in 2010 and trending toward a nearly monochromatic value by 2020.
WINTER observations continue to follow both sources through the 2024 and 2025 seasons, with both sources showing a continuous, gradual brightening in all three filters.
2.4 Archival imaging of the source environment
Both sources were observed by Spitzer/IRAC in all four channels as part of the GLIMPSE survey (Benjamin et al., 2003). We retrieved the data to provide both progenitor photometry and environmental context. Point-source photometry was obtained from the Spitzer source catalog (Spitzer Science, 2009), and wide-field color images were generated using archival IRAC mosaics from the CDS/P/SPITZER/color HiPS survey, accessed via the hips2fits service. We used the Aitoff projection centered on each FUor’s coordinates, with a field of view of 0.6 degrees and an output resolution of 1024 x 1024 pixels; the resulting FITS cubes were split into individual R, G, and B channels corresponding to the 8.0, 4.5, and 3.6 m IRAC bands, and combined into the composite image shown in Figure 1.
2.5 Spectroscopic observations
We obtained NIR spectroscopy of the sources with the SpeX spectrograph on the NASA Infrared Telescope Facility (Rayner et al., 2003) and the Folded-port Infrared Echellette (Simcoe et al., 2013, FIRE) on the Magellan telescopes (Figure 3). WNTR24-cua was observed with FIRE on UT 2024-07-04, while WNTR24-egv was observed with SpeX on UT 2022-10-08 and with FIRE on UT 2024-07-11. The SpeX data were taken in the SXD mode with a 0.8″slit, for a total exposure time of 30 minutes as part of program 2022B059 (PI: De). The data were reduced with the spextool pipeline (Cushing et al., 2004) while telluric and flux calibration were carried out using the xtellcor tool (Vacca et al., 2003). The FIRE data were taken in the prism cross-dispersed echelle mode with a 0.6” slit, which delivers spectra at R=6000 across m, and reduced with the PypeIt package (Prochaska et al., 2020).
3 Analysis
3.1 Environment
Both WNTR24-cua and WNTR24-egv lie next to nebulosity seen in the Spitzer images (Figure 1). WNTR24-cua is in the middle of some nebulosity but no obvious large HII region. Guo et al. (2024a) report a projected association between WNTR24-cua and a nearby star-forming region located at a heliocentric distance of 6.1 kpc (), based on the presence of a star-forming region within 5 arcmin of the source. They find, however, a significantly smaller kinematic distance of kpc (), derived from CO absorption line radial velocities. These values are not consistent within uncertainties, suggesting that WNTR24-cua likely lies in the foreground of the associated SFR. Guo et al. (2024a) caution that is only a rough estimate and may be affected by projection effects, while may be biased by unresolved binarity or peculiar motion. The offset in distance implies that WNTR24-cua may be forming in a more isolated or diffuse environment rather than within a clustered star-forming region.
WNTR24-egv lies in close proximity to the H ii region [DLW84] G35.2S, with G35.2N also nearby.. Using Gaia EDR3 parallax measurements ( mas; Gaia Collaboration 2020), we estimate a distance of kpc to the region. The projected angular separation between WNTR24-egv and the center of G35.2S is , corresponding to a physical offset of only pc at this distance, indicating a likely association. A dense molecular core (W48 89) identified by Keown et al. (2019) is essentially identical in position with WNTR24-egv. Many sources in the immediate vicinity were identified by Kuhn et al. (2021) as YSOs, although not the progenitor of WNTR24-egv.
Following the procedure outlined in Tran et al. 2024, we also estimate the distance to the outbursts by identifying nearby candidate YSOs with known distances. First, we take sources within 20’ of the FUors from the Spitzer (SEIP) source list (SSTSL2, Spitzer Science Center & Infrared Science Archive (IRSA)(2021), SSC) and the 2MASS All Sky Catalog of point sources (Cutri et al., 2003) and apply a color cut of and to identify sources with IR-color excess as likely YSOs. IR-excess contaminants, such as star-forming galaxies, are eliminated following the procedures outlined in Gutermuth et al. (2009).
This results in 466 and 273 candidate YSOs for WNTR24-cua and WNTR24-egv respectively. For WNTR24-cua, 17 of the candidate YSOs have counterparts in the Gaia DR3 catalog (Gaia Collaboration, 2020) with measured parallaxes, yielding a median distance of kpc. Due to the discrepancies in the estimated distances presented in Guo et al. 2024a, we use the estimate of 2.4 kpc for the analyses in Section 3.3. Similarly, for WNTR24-egv, 10 of the candidate YSOs have Gaia distances ranging from 0.2 to 2.2 kpc. However, their relatively uniform spatial distribution offers no clear basis for a reliable distance estimate, and many of the stars in this region are likely too obscured for Gaia detections. We therefore adopt the distance of the nearby H ii region at 2.52 kpc.
3.2 Spectroscopic features and classification
Both WNTR24-cua and WNTR24-egv are spectroscopically similar to other FUors at NIR wavelengths (as explored in detail in Connelley & Reipurth, 2018). The spectra are shown in Figure 3 with another FUor, Gaia17bpi (Hillenbrand et al., 2018), for comparison. All three spectra demonstrate a “mixed-temperature” spectrum characteristic of FUors. At the shortest infrared wavelengths (around 1 m), the spectra resemble a late-F to early-G type star at T6000 K. However, at longer wavelengths the spectra are consistent with an M-type star at T3000 K with deep molecular absorption features. For example, there is prominent absorption from the 12CO band-head (around ) and H2O band (m and m). Molecular absorption from VO and TiO is strongest in Gaia17bpi, present but weaker in WNTR24-cua, and not confidently detected in WNTR24-egv. Additionally, there are atomic lines from Al I, Ca I, Mg I, Na I, and Si I, along with H Paschen lines in both objects.
Both of the outbursts show strong He I 10830 blue-shifted absorption, which is evidence for a strong wind (Connelley & Reipurth, 2018). We estimate a FWHM velocity of approximately 231 km/s and 305 km/s for WNTR24-cua and WNTR24-egv, respectively, and corrected for instrumental broadening. Measuring against the blue end of the absorption feature, we estimate a maximum velocity of 221 km/s and 321 km/s, for WNTR24-cua and WNTR24-egv.
Additionally, both WINTER sources exhibit steep red continua across the near-infrared, consistent with heavily reddened, embedded systems. As shown in Figure 4, their spectra are significantly redder than Gaia 17bpi, suggesting a greater degree of circumstellar extinction and/or cooler outer disk temperatures. The continuum shape in both cases rises sharply from the to bands, with minimal flux at shorter wavelengths. This behavior is consistent with their non-detections in ground-based optical surveys: archival ZTF data show no detections in or bands down to a 5 limit of 20 mag, indicating that these outbursts are effectively invisible at optical wavelengths. These limits imply mag, consistent with their red continua. The detection of WNTR24-cua in Gaia is likely due to the broad Gaia bandpass, which extends red-ward to nearly 1 m and can capture flux in the far-red continuum where these objects still emit. The steep spectral slopes, in combination with their strong molecular absorption features, support the classification of these objects as deeply embedded FUor-type outbursts.
Based on these spectroscopic properties, we classify WNTR24-cua and WNTR24-egv as FUors following the criteria outlined in Connelley & Reipurth (2018). We measure the equivalent widths of CO, Na I, and Ca I, as outlined in Messineo et al. (2021), and compare to bona fide FUors from the Connelley & Reipurth (2018) sample (Figure 4). Both WNTR24-cua and WNTR24-egv fall within the bona fide sample, with notably large CO equivalent widths. Additionally, both sources exhibit the eight NIR spectral characteristics of FUors outlined in Connelley & Reipurth (2018). These characteristics include CO absorption and molecular band absorption, metallic, hydrogen, and He I lines, and changing of spectral type with wavelength.
3.3 Spectral energy distribution
Archival infrared imaging reveals that both WNTR24-cua and WNTR24-egv have luminous progenitors, enabling reconstruction of their pre- and post-outburst spectral energy distributions (SEDs; Figure 5) and investigating both the pre-main-sequence nature of the progenitors and the total luminosity increase during outburst. WNTR24-cua is detected in deep near- and mid-infrared surveys, including VISTA VVV Data Release 2 (Minniti et al., 2017), the Spitzer GLIMPSE Source Catalog (I, II, and 3D; Spitzer Science 2009), and DECaPS DR2. WNTR24-egv lacks optical detections but appears in 2MASS (source ID 18580350+0138499), Spitzer/GLIMPSE, the UKIDSS Galactic Plane Survey (Lucas et al., 2008), and WISE. WNTR24-cua was not reported in The Spitzer/IRAC Candidate YSO Catalog for the Inner Galactic Midplane (SPICY) catalog because the progenitor is too faint (Kuhn et al., 2021), while WNTR24-egv does have a listed progenitor in the catalog.
Figure 5 presents SEDs spanning the optical to mid-infrared for both sources before and after their outbursts. The pre-outburst SED for WNTR24-cua combines 2008 Spitzer and 2015 VISTA photometry, while the post-outburst data are from 2023 Gaia and 2024 WINTER and WISE data. For WNTR24-egv, pre-outburst data come from 2008 Spitzer and 2012 UKIDSS observations, with post-outburst photometry from 2024. Although these epochs are not contemporaneous, they represent comparable phases in each source’s evolution and are useful for characterizing the progenitor system.
We model the progenitors of each outburst with a combination of photospheric models (from Hauschildt et al., 1999) and black bodies approximating dust in the system. We fit the photospheric models to the - and -band data with a surface gravity of log(g)=4.0—typical for an M dwarf star—and a range of temperatures and extinctions. The addition of black body SEDs model the NIR and MIR excess which could result from an extended dust envelope around the progenitor. For WNTR24-cua, we find a range of models with K at extinctions of reasonably fit the data. The addition of a K black body helps model the MIR emission.
WNTR24-egv is similarly well modeled with a cool star, though slightly hotter and with higher extinction. The models span K with . For dust in WNTR24-egv, we use a two-component black body model with and K.
To place the SED-based extinction estimates in context, we also derive values from spectral dereddening and Galactic dust maps. First, we estimate the visual extinction () toward each source by dereddening their near-infrared spectra using the Cardelli et al. (1989) extinction law () until their continuum slopes match that of Gaia 17bpi, which has a known extinction of mag. This comparison yields relative extinctions of mag and mag, corresponding to values of mag for WNTR24-cua and mag for WNTR24-egv. We then compared these values to predictions from the Bayestar 3D Galactic dust map (Green et al., 2019). For WNTR24-cua (2.4 kpc), the cumulative Galactic extinction is mag, rising to only mag along the full line of sight, indicating that the source already lies behind most of the Galactic dust in this direction. In contrast, WNTR24-egv (2.52 kpc) shows mag out to its distance compared to mag for the entire column, implying that only part of the Galactic dust lies in the foreground and a substantial fraction remains behind the source. The much higher extinction estimate from the SED fit ( mag) therefore indicates significant local extinction in the circumstellar environment. For the analysis in Section 3.4, we adopt the extinction estimates from the SED fits.
The SEDs of the progenitor systems also enable classification of the proto-stars via their spectral indices. For data between 2.2 and 10 m, we get spectral indices of for WNTR24-cua and WNTR24-egv, respectively. Following the classification scheme in Greene et al. (1994), both sources are Class I YSOs. In other classification schemes extending to longer wavelengths, the progenitor of WNTR24-cua could be considered a “flat spectrum” type YSO, similar to the FUor discussed in Tran et al. (2024). These spectral indices together with the dust excess and high extinction indicate the objects are Class I, although SPICY reports WNTR24-egv as a Class II object based on [4.5] - [8.0] colors (Kuhn et al., 2021).
3.4 Luminosities and mass accretion rates
We estimate the mass accretion rates of both FUor outbursts by combining physical parameters derived from pre-main-sequence stellar models with bolometric luminosities obtained from the SEDs. For each system, we integrate the extinction-corrected SED between 1 and 30 m to determine the total luminosity before and after outburst (following Gordon et al., 2023). The post-outburst photometry used in this analysis was obtained in 2024, years after the outburst onset for each event.
For WNTR24-cua, we adopt mag and find a pre-outburst luminosity of and post-outburst luminosity of . In WNTR24-egv, where mag, we estimate and .
To estimate the stellar mass and radius of the progenitors, we interpolate the 1 Myr pre-main-sequence evolutionary models of Baraffe et al. (2015) at the effective temperature and pre-outburst bolometric luminosity of each system. These luminosities include contributions from both the stellar photosphere and circumstellar dust re-emission. Thus, the inferred stellar masses and radii should be regarded as upper limits; the true stellar values may be lower if a significant fraction of the bolometric luminosity arises from non-stellar components. For WNTR24-cua, adopting K and , we infer and . For WNTR24-egv, with K and , we find and .
Assuming the excess luminosity is powered by accretion from a circumstellar disk, we use the standard expression for accretion luminosity,
with as the adopted inner disk truncation radius (Gullbring et al., 1998). Applying this relation, we find that in WNTR24-cua, measured approximately 7 years after the onset of its outburst, corresponds to an accretion rate of . For WNTR24-egv, the larger increase of , derived from observations taken roughly 10 years after the outburst began, implies .
These estimates fall within the expected range for FUor outbursts. Typical accretion rates of – have been reported both in observational surveys (e.g., Audard et al., 2014) and recent theoretical studies (e.g., Nayakshin et al., 2024), placing WNTR24-cua at the lower end of the distribution and WNTR24-egv near the median. The modest accretion rate inferred for WNTR24-cua is consistent with a low-mass progenitor and is similar to that of Gaia 17bpi, a low-luminosity FUor with comparable properties.
Given the uncertainty in the distance to WNTR24-cua, which ranges from kpc (calculated in Section 3.1) to the two proposed values of and kpc from Guo et al. (2024a), we scale the luminosities and accretion rates accordingly. At 3.0 kpc, the post-outburst luminosity increases to and the inferred accretion rate to . At 6.1 kpc, these values become and . These values remain within the expected range for FUors but suggest that WNTR24-cua may be more luminous than implied under the nominal 2.4 kpc assumption.
4 Discussion and Summary
WINTER identified two new FUor outbursts that are bright in the near-infrared, also detected in the mid-infrared, but largely obscured at optical wavelengths. WNTR24-cua is a previously characterized FUor and WNTR24-egv is a newly classified FUor from this sample. Long-term monitoring with surveys like WISE and Gaia show the rise and color evolution of the bursts, revealing years-long events with slow declines. WINTER observations of the events confirmed that these objects were bright in the NIR and prompted rapid spectroscopic follow-up. The spectra for both events show cool objects with mixed spectral types with evidence of winds or outflows. Furthermore, the spectra confirm the events as FUors in the Connelley & Reipurth 2018 classification scheme. Compared to the larger sample of FUors, the two WINTER events show notably high CO equivalent widths (Figure 4) and steep spectral slopes.
Although WNTR24-cua was previously reported, we present its Gaia, WINTER, and NEOWISE light curve and progenitor properties. This indicates that the mid-infrared rise seen in NEOWISE started shortly before (1 year) the Gaia optical outburst. This is very similar to the object in Guo et al. 2024b, which also shows a later time re-brightening as seen in this object.
This paper presents the spectroscopic confirmation of WNTR24-egv as a bona fide FUor, previously identified as a candidate based on its mid-infrared variability. Despite the lack of any optical signature during the outburst, NEOWISE photometry reveals that the event began around 2014 and has continued to brighten steadily in the infrared. Given the estimated distance and the high extinction inferred from SED fitting—well in excess of the expected foreground extinction—the source is likely a deeply embedded YSO. This level of obscuration naturally explains the absence of any optical counterpart throughout the duration of the outburst.
Though both objects exhibit spectroscopic similarity, their outburst timescales differ markedly. WNTR24-cua shows a rapid rise, with the infrared brightening (seen in WISE and VVV ) about a year before the optical Gaia outburst (Figure 2). In contrast, WNTR24-egv is undergoing a slower outburst, with longer wavelengths likely leading shorter ones, as evidenced by W2 rising more gradually and persisting longer than W1. The color evolution of both sources aligns with outside-in propagation models, which predict the infrared to brighten before the optical. Cleaver et al. (2023) show that the delay between infrared and optical rise correlates with accretion rate, with more luminous (i.e., higher accretion rate) events exhibiting longer delays. It is therefore possible that WNTR24-egv has not yet reached its optical peak, and the current infrared brightening is an early precursor—as some models predict the IR outburst can precede the optical by decades (e.g., Cleaver et al., 2023).
Overall, these studies highlight the importance of multi-color monitoring of the Galactic plane to understand the triggering mechanisms of FUors. The classical thermal instability model predicts that the IR rise lags behind the optical (Bell et al., 1995). In contrast, magnetorotational instability (MRI) models predict an infrared precursor, with the delay depending on the luminosity and mass accretion rate. For the largest outbursts—such as FU Orionis itself—Cleaver et al. (2023) argue that the outburst originates far out in the disk, producing an infrared precursor decades before the optical rise. However, the delay is shorter for smaller peak outbursts, where the instability occurs closer to the star. Models by Nayakshin & Elbakyan (2024) instead suggest that the evaporation of a close-in planet can trigger FUor outbursts; in this case, the delay does not depend on the mass accretion rate. Guo et al. (2024b) find that such a scenario is more consistent with similar delays observed in outbursts with varying peak accretion rates.
While models disagree on whether the IR–optical delay should scale with accretion rate, distinguishing between these scenarios requires monitoring across multiple wavelengths and cadences from weeks to decades. Prior infrared surveys like VVV and PGIR lacked the multi-color, high-cadence coverage needed to resolve early outburst evolution, while NEOWISE operated only at mid-infrared wavelengths where spectral energy distributions are less sensitive to changes in temperature. Optical time-domain surveys are largely blind to these events due to extinction. WINTER’s near-infrared capabilities—combining , , and band coverage with regular cadence—directly address this gap, offering sensitivity to the color evolution that discriminates between theoretical models (e.g., inside-out versus outside-in instability propagation; Bell et al. 1995; Cleaver et al. 2023). Although WINTER observations of the two FUors studied here began only after their outbursts were underway, future real-time detections will enable systematic tests of these models.
When the Galactic plane is visible, WINTER continues its survey of the plane and nearby star-forming regions, enabling ongoing monitoring of the two FUor events presented here, as well as the discovery of new outbursts. The survey operates on a two-week cadence in - and -bands, targeting northern fields including active star-forming regions such as W40 and the North America Nebula. Near-IR coverage will catch obscured events missed by optical surveys and help fill in an infrared gap now that mid-IR monitoring with WISE has concluded. In particular, WINTER is prioritizing coverage of the northern fields selected for the upcoming Roman Galactic Plane Survey, providing long-term near-infrared monitoring ahead of Roman’s deeper surveys. Roman observations will also provide resolved follow-up of outbursts seen in WINTER.
The 2-week cadence and time-domain infrastructure of WINTER enable rapid spectroscopic follow-up of newly detected outbursts, allowing classification of short-duration bursts or while long-duration events are still evolving (Figure 6). Early spectroscopy is particularly valuable for capturing transient features during the rise phase, when the structure of the accretion disk is still forming. Timely classification also helps distinguish between different types of transients—such as FUors, EXors, classical novae, or planetary engulfment events (Audard et al., 2014; Chomiuk et al., 2021; De et al., 2023). In addition, the high cadence improves sensitivity to short-duration, low-amplitude events that are easily missed by slower surveys. WINTER is uniquely positioned to detect and spectroscopically characterize both short-lived transients and long-duration FUor-type events.
As both NEOWISE and VVV have now concluded operations, WINTER will provide ongoing infrared coverage of the Galactic plane in the years leading up to the launch of the Roman Space Telescope and its planned Galactic plane survey (RGPS) (Sanderson et al., 2024). WINTER’s high-cadence, multi-band near-infrared monitoring complements the deep optical observations from the Vera C. Rubin Observatory (Ivezić et al., 2019), the high-resolution infrared imaging expected from Roman, and deep single-band NIR monitoring from the future Cryoscope survey (Kasliwal et al., 2025). Together, these facilities will enable more complete identification and characterization of YSO outbursts, spanning a wide range of evolutionary stages and environments. In particular, WINTER’s early detections will ensure that FUors erupting during or after the RGPS can be linked to Roman photometry of their faint progenitors—for example, the progenitor of WNTR24-cua was barely detected in pre-outburst imaging. Additionally, for active FUors, Roman will also offer the opportunity to study spatially resolved outflows, complementing what has previously only been achievable through targeted Hubble Space Telescope follow-up. As WINTER continues its Galactic plane survey, it is building a growing sample of eruptive YSOs—including both classical FUors and more rapidly evolving events—that will serve as a valuable resource for multi-wavelength studies of episodic accretion in the time-domain era.
5 Acknowledgments
WINTER’s construction is made possible by the National Science Foundation under MRI grant number AST-1828470 with early operations supported by AST-1828470. Significant support for WINTER also comes from the California Institute of Technology, the Caltech Optical Observatories, the Bruno Rossi Fund of the MIT Kavli Institute for Astrophysics and Space Research, the David and Lucille Packard Foundation, and the MIT Department of Physics and School of Science. D.F.’s contribution to this material is based upon work supported by the National Science Foundation under Award No. AST-2401779. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. Furthermore, we acknowledge the support of the National Aeronautics and Space Administration through ADAP grant number 80NSSC24K0663. This work is based on observations obtained at the Infrared Telescope Facility, which is operated by the University of Hawai‘i under contract 80HQTR19D0030 with the National Aeronautics and Space Administration. We acknowledge ESA Gaia, DPAC and the Photometric Science Alerts Team (http://gsaweb.ast.cam.ac.uk/alerts). This research made use of hips2fits,333https://alasky.cds.unistra.fr/hips-image-services/hips2fits a service provided by CDS.